### v1

# Nonlinear dynamics of confined cell migration—modeling and inference

Abstract
The motility of eukaryotic cells is strongly influenced by their environment, with confined cells often developing qualitatively different motility patterns from those migrating on simple two-dimensional substrates. Recent experiments, coupled with data-driven methods to extract a cell’s equation of motion, showed that cancerous MDA-MB-231 cells persistently hop in a limit cycle when placed on two-state adhesive micropatterns (two large squares connected by a narrow bridge), while they remain stationary on average in rectangular confinements. In contrast, healthy MCF10A cells migrating on the two-state micropattern are bistable, i.e., they settle into either basin on average with only noise-induced hops between the two states. We can capture all these behaviors with a single computational phase field model of a crawling cell, under the assumption that contact with non-adhesive substrate inhibits the cell front. Our model predicts that larger and softer cells are more likely to persistently hop, while smaller and stiffer cells are more likely to be bistable. Other key factors controlling cell migration are the frequency of protrusions and their magnitude of noise. Our results show that relatively simple assumptions about how cells sense their geometry can explain a wide variety of different cell behaviors, and show the power of data-driven approaches to characterize both experiment and simulation.
I Introduction
Living cells navigate intricate spatial constraints and complex environments and respond to them by changing their motility and behavior [1, 2]. This can range from developing biased migration in response to asymmetric micropatterns [3, 4, 5] to altering mechanisms of migration [6, 7, 8] or developing completely different migration patterns such as rotations and oscillations [9, 10, 11, 12]. How can we characterize all of these complex types of migration and determine how they arise from the influence of confinement? One promising recent approach is the use of data-driven models [13, 14, 15, 16, 17, 18], in which equations of motion are learned directly from experimental data. In particular, we focus on a recent study on the migration of cancerous MDA-MB-231 and healthy MCF10A cells [14], which summarized experimental observations into a data-driven representation of the dynamics of the cell position $x$ and velocity $v$. Using this representation, cells were characterized with the language of nonlinear dynamics. For instance, cancerous MDA-MB-231 cells were well described by a limit cycle in $x$ and $v$ when placed on a “two-state geometry”, consisting of two square basins connected by a narrow bridge (Fig. 1). This limit cycle characterizes cells persistently hopping from basin to basin. This is strongly dependent on the presence of the narrow constriction: when MDA-MB-231 were placed on a rectangular substrate, they remained stationary on average. By contrast, MCF10A cells were bistable on the two-state micropattern, settling into either basin on average and only executing noise-induced hops [14].
Why does the two-state geometry break the symmetric migration of cancerous cells and cause them to limit-cycle? Why do MDA-MB-231 and MCF10A cells behave differently? To what extent does the experimental measurement of these nonlinear dynamics constrain more complicated models of cell motility? We attempt to construct a single model of a cell that can capture the rich landscape of behaviors observed by [14]. We use the phase field framework to describe our cell motion, with a simple assumption that a cell’s polarity is inhibited when the cell leaves the micropattern. We apply the data-driven approach of [14] to determine the equation of motion for the cell’s velocity and position. We show that our model can successfully recapitulate limit cycle, bistable, and stationary behaviors depending on cell attributes and geometry of the micropattern. Our model suggests that larger and softer cells tend to limit-cycle on two-state micropatterns, while smaller and stiffer ones are more likely to be bistable. We also study the role of other perturbations of our cell polarity model, and characterize the variability in the predictions of the data-driven model due to finite sampling. Our results suggest that the qualitative difference between bistable and limit cycle behavior can be achieved by relatively small quantitative changes in mechanical, physical, or polarity properties of the cell, including cell size and cell tension. These results highlight the power of data-driven approaches to detect fine changes in cell motility.
II Model
We model a single cell within the phase field framework [19, 20, 21, 22, 9, 23], which describes objects with arbitrary and continuously deforming shapes. Because the experiments of [14] confine the cell on two-dimensional adhesive micropatterns, we simplify our model to two dimensions and adopt a “top view” of the cell (see Fig. 3). The cell shape is set by a field $\phi({\bm{r}},t)$, which is zero outside the cell and one inside it, implicitly defining the boundary at $\phi=1/2$. The evolution of the cell boundary is governed by energy minimization and advection of the interface [20, 24, 25],
$\frac{\partial\phi({\bm{r}},t)}{\partial t}+{\bm{v}}({\bm{r}},t)\cdot\nabla\phi=-\mathcal{M}\ \frac{\delta\mathcal{F}}{\delta\phi},$ (1)
where $\mathcal{M}$ is the transport coefficient, and ${\bm{v}}({\bm{r}},t)$ is the velocity field of the cell driving its interface. The free energy is $\mathcal{F}=\mathcal{F}_{\mathrm{CH}}+\mathcal{F}_{\mathrm{area}}$. The Cahn-Hilliard energy is [20]
[…]
which is minimal at $\phi=0,1$ and penalizes deformations of the interface. $\lambda$ has units of length and sets the phase field interfacial thickness, while in two dimensions, $\gamma$ sets the line tension of the cell with units of energy/length. The term $\mathcal{F}_{\textrm{area}}$ models cell incompressibility by penalizing deviations of cell area from its target $\pi R_{0}^{2}$ [20],
[…]
In the absence of cell motility, these energy terms equilibrate the cell to a circle with radius $R_{0}$.
We introduce cell motility through the velocity field $\bm{v}(\bm{r},t)$. Modeling the cell as an overdamped system, we locally balance forces per unit area and write [20]
$\eta{\bm{v}}=\frac{\delta\mathcal{F}}{\delta\phi}\nabla\phi+{\bm{f}}_{\mathrm{motility}}.$ (4)
The left hand side represents the friction force per unit area, while the first term on the right denotes the velocity of the cell boundary as governed by energy minimization. The last term is the active motility force per unit area, which depends on cell polarity and drives the cell front.
II.1 Micropattern geometries
The experiments of [14] use an adhesive fibronectin micropattern to effectively confine the cells—cells can only adhere to the substrate on the micropattern. We incorporate these micropatterns in our phase field framework using a static field $\chi(\bm{r})$. We set $\chi$ to be zero in the interior and one in the exterior of the micropattern, transitioning smoothly at the boundary (Fig. 1). We assume that the cell interacts with the substrate through its polarity, which controls cell motility. This neglects any adhesive energy to the micropattern [26, 21].
Figure 1: A representation of the two substrate geometries by a phase field $\chi(\bm{r})$, which is 0 (black) in the interior and 1 (white) in the exterior. The two-state geometry has two basins measuring $38{\mu\mathrm{m}}\times 38{\mu\mathrm{m}}$ that are connected by a $10{\mu\mathrm{m}}\times 35{\mu\mathrm{m}}$ bridge. The rectangular substrate measures $111{\mu\mathrm{m}}\times 38{\mu\mathrm{m}}$.
II.2 Defining cell polarity and motility
Migrating cells have an underlying chemical polarization, indicating the areas of the cell that are likely to protrude (“front-like”) and those likely to contract (“rear-like”) [27]. This can include asymmetric distribution of Rho GTPases, with Rac1 activity driving the cell front through lamellipodial extensions, and RhoA promoting myosin contractility in the rear [28]. Rather than explicitly modeling the dynamics of one or more Rho GTPases [22, 9, 29, 30, 31, 32], we summarize cell polarity with a single spatiotemporal field $\mathbb{P}({\bm{r}},t)$ whose sign determines whether the cell boundary protrudes ($\mathbb{P}>0$) or contracts ($\mathbb{P}<0$). This leads us to write the motility force as
$\bm{f}_{\mathrm{motility}}=\alpha\mathbb{P}(\bm{r},t)\phi^{2}(1-\phi)^{2}(1-\chi)\bm{\hat{n}}.$ (5)
Here, $\alpha$ is a constant, $\bm{\hat{n}}\equiv-\nabla\phi/|\phi|$ is the unit vector normal to the cell interface, and the phase field terms ensure that motility forces exist only where the cell interface (i.e. $\phi^{2}(1-\phi)^{2}>0$) is in contact with the micropattern (i.e. $01-\chi>$ 0)[28, 21, 33].
There is no single established approach for modeling the complex dynamics of cell polarity [34]. Here, we evolve $\mathbb{P}(\bm{r},t)$ using stochastic fluctuations in polarity at random locations on the cell boundary, inhibition of polarity in the absence of the micropattern, inhibition of polarity when the cell’s tension surpasses a certain threshold, and a basal decay of polarity. Mathematically, we write
$\begin{split}\frac{\partial\mathbb{P}(\bm{r},t)}{\partial t}&=\beta(t)f\left(|\bm{r}-\bm{\theta}(t)|;\sigma\right)\delta(t\text{ mod }\tau_{f})\phi\\ &-\frac{1}{\tau_{\chi}}\chi\phi\\ &-\frac{1}{\tau_{s}}\left(\frac{s}{s_{0}}\right)^{2}H(s-s_{0})\\ &-\frac{1}{\tau}\mathbb{P}(\bm{r},t).\end{split}$ (6)
The first term introduces radially-symmetric patches $f\left(|\bm{r}-\bm{\theta}(t)|;\sigma\right)=e^{-|\bm{r}-\bm{\theta}(t)|^{2}/2\sigma^{2}}$ that activate polarity at locations $\bm{\theta}$ on the cell boundary. The amplitude of the patch is a random variable $\beta\sim\mathrm{Gaussian}(\mu_{\beta},\sigma_{\beta})$. We choose $\mu_{\beta}>0$ so these patches locally increase the polarity on average, but $\beta(t)$ can experience large fluctuations, including negative values. Lastly, we write $\delta(t\text{ mod }\tau_{f})\phi$, where $\delta(t)$ is the Dirac delta, which lets us introduce these noisy activity patches at a controlled frequency $\tau_{f}^{-1}$ and ensure they are confined within the cell interior. To complete this description, we must specify a location $\bm{\theta}$ for the polarity patch every $\tau_{f}$. The polarity only drives protrusion and retraction at the interface, so we choose $\bm{\theta}$ to be a contour point along the boundary of the cell ($\phi=1/2$). We assume cells choose these locations with two important factors: 1) a generic positive feedback, where existing polarity promotes new polarity, akin to other polarity models [27, 35], and 2) feedback from successful adhesion, where new cell-substrate adhesion promotes polarity [36]. We choose the locations of polarity patches according to these principles by sampling $\bm{\theta}$ from a probability $p(\bm{\theta})$ every $\tau_{f}$, where
$p(\bm{\theta})\sim\left[1-\chi(\bm{\theta}^{\prime})\right]\times\mathbb{P}(\bm{\theta}),$ (7)
and $\bm{\theta^{\prime}}=\bm{\theta}+\ell\bm{\hat{n}}(\bm{\theta})$ is a point a distance $\ell$ normally outward from the contour (Fig. 2). In the $\left[1-\chi(\bm{\theta}^{\prime})\right]$ term, the cell is essentially measuring to discover if the adhesive micropattern is present (indicated by regions where $01-\chi>$ 0) at the point $\bm{\theta}^{\prime}$, and if so, enhancing polarity at $\bm{\theta}$. The idea behind this “filopodial probability” is that small stochastic protrusions (filopodia) that are not explicitly represented in our model of the cell boundary test the region near the cell edge, and upon contact with the adhesive micropattern, mature and reinforce polarity [5, 28]. Lastly, our assumption that the probability to pick a contour point $p(\bm{\theta})$ is proportional to the polarity at that contour point $\mathbb{P}(\bm{\theta})$ is a simple positive feedback mechanism, where higher polarity at $\bm{\theta}$ means higher likelihood of sampling it for subsequent addition of an activity patch. Taken together, sampling from the probability of Eq. (7) generates more activity patches at points that have a reasonable amount of substrate in front of them and currently higher polarity values (Fig. 2).
The second term in Eq. (6) says that cells decrease their polarity in regions where they lose contact with the micropattern (places where $\chi\phi\neq 0$). This term, along with the filopodial probability in Eq. 7, are the only two mechanisms by which the cell interacts with the micropattern.
The third term in Eq. (6) uniformly inhibits polarity when the cell perimeter $s$ exceeds a threshold $s_{0}$ (indicated by the Heaviside function $H(s-s_{0})$). This assumption is motivated by experiments where membrane tension was observed acting as a global, long-range inhibitor of actin polymerization in neutrophils [37] (also see [38, 39]).
The last term in Eq. (6) decays the polarity over a timescale $\tau$, setting the characteristic time over which the cell can maintain a given polarity.
Figure 2: A cartoon of how polarity at contour points and their proximity to the micropattern edge combine to define the probability $p(\bm{\theta})$. The presence of the micropattern is measured at $\bm{\theta}^{\prime}$, which is obtained by projecting each contour point normally outward a distance $\ell$ (red dots). In this sketch, the highest probability is associated with the contour point that measures the highest polarity and amount of micropattern ahead of it.
II.3 Parameter setting
The numerical values listed in Table 1 define our default simulation cell, which is calibrated to behave similarly to cancerous MDA-MB-231 cells in [14]. We take the values of $\lambda$, $\kappa$, $\eta$, and $\mathcal{M}$ from our earlier work [26], where we calibrated to a different experimental study of MDCK cells migrating on narrow channels [40]. As in our earlier work [26], we define many of our parameters relative to an arbitrary line tension scale $\gamma_{0}$—the results are not sensitive to the true value of the line tension, only how these different parameters relate to $\gamma_{0}$.
Parameter
Value
Units
$\gamma$
0.8
$\gamma_{0}$
$\lambda$
4.8
${\mu\mathrm{m}}$
$\kappa$
6000
$\gamma_{0}\ {\mu\mathrm{m}}$
$R_{0}$
18
${\mu\mathrm{m}}$
$\eta$
0.67
$\gamma_{0}\ \rm min\ {\mu\mathrm{m}}^{-3}$
$\mathcal{M}$
0.75
${\mu\mathrm{m}}\ (\gamma_{0}\ \rm min)^{-1}$
$\mu_{\beta}$
0.94
$\rm min^{-1}$
$\sigma_{\beta}$
4.7
$\rm min^{-1}$
$\sigma$
6.7
${\mu\mathrm{m}}$
$s_{0}$
$2\pi(1.5R_{0})$
${\mu\mathrm{m}}$
$\tau_{f}$
2.2
$\rm s$
$\tau_{\chi}$
9.6
$\rm s$
$\tau_{s}$
8
$\rm min$
$\tau$
4
$\rm min$
$\ell$
6
${\mu\mathrm{m}}$
$\alpha$
1.4
$\gamma_{0}\ {\mu\mathrm{m}}^{-2}$
Table 1: A list of simulation parameters along with the values that define our default cell, which is calibrated to behave similarly to MDA-MB-231 cells in [14]. See Appendix A for parameters relevant to the numerical integration scheme.
The constraint penalizing deviations from the preferred area, $\kappa$, is set relatively strongly to ensure the area of the default cell varies less than $0.5\%$ from its target. The timescale over which cell polarity decays to zero in regions where the cell is near substrate edges, $\tau_{\chi}$, is set so that when the cell boundary leaves the micropattern, the polarity rapidly becomes negative, leading to the membrane contracting away from the micropattern edge, preventing the cell from leaving the pattern entirely. The perimeter-induced degradation in polarity prevents the cell from becoming too elongated. Since we do observe that cells in [14] can be quite elongated, we cannot set the timescale of this process $\tau_{s}$ too short, or it will prevent cell stretching. We choose $\tau_{s}$ to be on the order of the global decay $\tau$, and note that the decay rate will be faster for larger values of cell perimeter. We set the magnitude of the motility force $\alpha$ so that our simulated cells achieve speeds and accelerations that roughly match the experiments. The rest of the parameters in Table 1 will be discussed in §III as we consider their variations and effect on cell behavior.
II.4 Inferring motility patterns from simulations
We begin each simulation by either positioning the cell randomly in one of the two basins of the two-state substrate or uniformly distributing it across the rectangular micropattern. Upon initialization, we numerically evolve the phase field equations of motion for 48 hours (see Appendix A). Unlike the experiments, which recorded the position of cell nucleus, we track the $x$ position of cell center of mass. While nuclear dynamics could be modeled with another phase field [9], we neglect it here for simplicity and claim that cell center of mass is a fair approximation of cell nucleus for most of the morphologies we observe.
We adapt the approach of Brückner et al. [14] to our simulations by representing the dynamics of the cell center-of-mass position and velocity with the stochastic equation of motion,
$\dot{v}=F(x,v)+\sigma(x,v)\eta(t),$ (8)
where $\eta(t)$ is a Gaussian white noise with $\langle\eta(t)\rangle=0$ and $\langle\eta(t)\eta(t^{\prime})\rangle=\delta(t-t^{\prime})$. Here, we focus only on $F(x,v)$, which denotes the average center-of-mass acceleration of the cell when it has position $x$ and velocity $v$. Since cell motility is stochastic, we perform $N=4800$ simulations for every cell-micropattern pair. Mirroring the approach of [14], we compute velocities and accelerations as numerical derivatives, construct the stochastic acceleration field $\dot{v}(x,v)$ from each simulation, and define the deterministic footprint by computing the average acceleration $\dot{v}$ conditional on the value of $x$ and $v$, i.e. $F(x,v)\equiv\langle\dot{v}|x,v\rangle$ (refer to Appendix B for details on this inference scheme).
Similar to [14], we interpret $F(x,v)$ as the blueprint that defines the deterministic migration of the cell on a given micropattern geometry. One way to visualize the motility pattern encoded by $F$ is to trace its streamlines—trajectories that outline the instantaneous direction of motion in phase-space. For instance, we can begin with a point in the $xv$ phase-space, say $(x_{0},v_{0})$, and integrate it one timestep forward to ($x_{0}+v_{0}dt,\ v_{0}+dtF(x_{0},v_{0})$). Repeating this procedure will eventually reveal the phase-space trajectory, i.e., the streamline, dictated by $F(x,v)$ when starting at $(x_{0},v_{0})$. A streamline will terminate once it reaches a point at which both acceleration and speed are zero, i.e., the point $(x_{f},v_{f})$ such that $F(x_{f},v_{f})=0$ and $v_{f}=0$.
We generally see streamline plots that have one of three typical forms: 1) a single limit cycle, where streamlines converge onto a closed loop, 2) “stationary” behavior, where streamlines all converge to $v=0$ but do not have a strong dependence on $x$, and 3) bistability, where two sets of disconnected streamlines converge to zero speed at either basin $x=\pm x_{\mathrm{basin}}$. When a streamline terminates, we mark its end point with an $\bm{\times}$ and interpret it as the point to which the cell would relax to if its equation of motion were to evolve without any noise. When multiple terminal points occur, we group trajectories with common endpoints and represent each group with distinct colors. We then draw the corresponding separatrices with solid lines to delineate the boundaries between regions of phase space that are attracted to different terminal points.
The blueprints $F(x,v)$ are inferred from a finite number of observations, leading to some uncertainty in the values of $F(x,v)$. This can lead to notable uncertainty in some terminal points and separatrices. We use a bootstrapping technique to show the scale of expected variations in our streamline plots, terminal points, and separatrices in Appendix E.
III Results
III.1 Model predicts limit-cycles on two-state and stationary behavior on rectangular micropatterns
Can our proposed model of the cell simultaneously capture limit-cycles on two-state and stationary behavior on rectangular geometries, as observed for cancerous MDA-MB-231 cells by [14]?
Figure 3 plots one hopping cycle of the cell as it navigates the two-state geometry in our simulations (see Movie M1 for a longer trajectory). The cell adopts a singly-polarized morphology with a clear cell front that is rich in polarity and elongates to travel across the bridge between the two square micropatterns. We plot trajectories of the cell center of mass from a few simulations in Fig. 3. Our default cell, calibrated to MDA-MB-231 cells, is persistently hopping between basins, as MDA-MB-231 cells in [14] do. We then infer the deterministic acceleration footprint $F(x,v)$, which is a concise way of summarizing the migration of the cell observed across 4800 simulations. Figure 3 plots this footprint in the $xv$ space using a colormap. When we compute the streamlines (as described in Section II.4), we find that the motion of the cell described by $F(x,v)$ forms a closed loop in the $xv$ space. This mean that our model, at its default parameters, creates a limit-cycle behavior, where cells hop persistently between basins. This finding is consistent with the limit cycle behavior reported in [14], which we have reprinted in Fig. 3.
Our default cell behaves entirely differently on the rectangular geometry. Figure 3 plots a few snapshots of the cell as it migrates on the rectangular substrate, where we observe the cell adopting a bipolar shape and undergoing a “tug-of-war” with competition between two protrusions, similar to that of MDA-MB-231 cells in experiments [14]. The center of mass trajectories in Figure 3 show that the cell is not persistently migrating but has small fluctuations and occasional larger steps. Note that unlike migration in the two-state geometry, cells can remain near $x=0$ on the rectangular micropattern for long periods of time, not just transiently while hopping. Mirroring the above analysis, we infer $F(x,v)$ from simulations and plot it in Figure 3. This time, tracing the streamlines of $F(x,v)$ reveals trajectories that bring the cell quickly to the $v=0$ line, with termination points distributed along $x$, suggesting that our model on the rectangular geometry, at its default parameters, generates an $F(x,v)$ where cells quickly relax to become stationary with a position near its initial position. We call this sort of streamline plot “stationary” throughout the paper. This finding is consistent with the similar stationary behavior of cancerous MDA-MB-231 cells observed in [14], which we have reprinted in 3.
As a side note, we should mention that the specific locations of the terminal points in Fig. 3 are somewhat uninformative. For this stationary case, the cell will have near-zero $F(x,v)$ near $v=0$ for a broad range of values of $x$, small sampling error can change the positions of the $\times$-labeled termination points. In later plots that show this “stationary” dynamics, we do not label putative termination points.
Case 1: Default cell limit-cycles on two-state geometry
Case 2: Default cell is stationary on rectangular geometry
Case 3: Decreasing cell size ($R_{0}=15{\mu\mathrm{m}}$) causes bistability on two-state geometry
Figure 3: Overview of limit-cycle (3-3), stationary (3-3), and bistable (3-3) behavior. Each case shows sample snapshots of typical morphologies and polarity (red: high, blue: low) (3, 3, 3), outlines center-of-mass trajectories (3, 3, 3), plots $F(x,v)$ inferred from simulations (3, 3, 3), and reprints the experimental footprints of [14] with permission for comparison (3, 3, 3). Streamlines are colored by equilibrium point; $\times$ indicates equilibrium points where streamlines terminate. Solid black lines are separatrices—crossing this line changes the long-run behavior of the cell. See Section II.4 for details.
III.1.1 Geometry sensing and the origin of limit-cycles
Let’s consider migration in the two-state geometry. When the cell is inside a basin and large enough to fill it up, most of its interface lies near the edges of the confinement except the part facing the bridge. Since we degrade polarity in regions where the cell overlaps with the micropattern (Eq. (6)), and suppress the generation of activity in regions that are near to the end of the micropattern (filopodial probability of Eq. (7)), sustained polarity can only develop in regions facing the bridge. This region further experiences a targeted increase in average cell polarity due to the positive feedback assumption of Eq. (7). The tendency to polarize toward free micropattern space and positive feedback work cooperatively to yield a persistent, singly-polarized cell that migrates through the bridge and into the other basin. Once on the other side, the cell reverses in polarity as the bridge now faces the opposite direction and repeats the cycle above. The result is a persistently polarized cell that limit-cycles on the two-state geometry. Interestingly, we find that the filopodial probability of Eq. (7) is dispensable—the assumption that leaving the micropattern decreases cell polarity is sufficient for reproducing limit-cycles in the two-state micropattern (see Section III.6 for details).
III.1.2 Origin of stationary behavior on rectangular substrates
Let’s now consider the evolution of cell polarity on the rectangular substrate. Suppose polarity is initially constant in space. Since the cell is confined at its top and bottom, sustained polarity can only develop at the sides, yielding elongation in ±$\pm x$. Given the stochastic evolution of cell polarity and our assumption of positive feedback, which favors a one-side-wins-all configuration, we would initially expect the cell to be singly-polarized and have one clear front, just like it is on the two-state substrate. Instead, we observe the average cell shape is bipolar (see Fig. 3 and Movie M2). This bipolarity arises because of a coupling between the cell polarity and the evolution of cell shape. When we numerically solve the polarity model of Eqs. (6) and (7) on a static phase field (see Appendix C), we consistently arrive at singly-polarized cells. What is different about a cell whose boundary dynamically moves? The polarity field $\mathbb{P}(\mathbf{r},t)$ is confined within the body of the cell, and while it responds to motion of the cell boundary, it does not do so instantaneously. When a high-amplitude patch of activity hits the contour ${\bm{\theta}}_{k}$, it increases the polarity in that region and leads to a protrusion of the cell membrane. This protrusion moves ${\bm{\theta}}_{k}$ normally outward and away from its previous location. Because $\mathbb{P}$ does not instantly migrate with the cell boundary, the polarity measured at the same contour point shortly after protrusion is not nearly as high as before—i.e. polarity can be implicitly suppressed by protrusion. We hypothesize that this negative feedback can prevent our positive feedback terms in Eq. (7) from leading to a singly-polarized cell. Our discovery that cell shape dynamics are necessary here is part of a broader realization that complex behaviors can emerge from the coupling of cell polarization to cell shape [41, 24, 42].
III.2 Cell size modulates bistability on two-state substrates
Thus far, we have demonstrated our model’s ability to recapitulate limit cycles on two-state and stationary behavior on rectangular substrates. Now, we ask which properties of the cell might dictate whether it shows limit cycles or bistability on the two-state micropattern. These are potential candidates for being key differences between MDA-MB-231 cells and MCF-10A cells. In this section, we focus on perturbing cell size, which we control through the target radius $R_{0}$, leaving all other parameters the same as those in Table 1.
Figure 3 presents a few simulation snapshots of a smaller cell with $R_{0}=15{\mu\mathrm{m}}$ on the two-state geometry. According to these snapshots, the smaller cell experiences difficulty hopping between the basins and appears more likely to repolarize back into either state of the micropattern. Note that while the snapshots we highlighted don’t include a hopping event, smaller cells can indeed hop between the basins of the micropattern, and when they do, their morphology is highly similar to that of the default cell in Fig. 3 (see Movie M3 for a simulation that includes hopping). The lack of persistent hopping in smaller cells is readily apparent from the center-of-mass trajectories of Figure 3, where we observe smaller cells to remain in a given basin for much longer times with infrequent hopping to the other side. When we infer $F(x,v)$ and trace its streamlines, we find, in contrast with the limit-cycle dynamics of Fig. 3, trajectories that terminate either at the left or right basin. We color trajectories that terminate in the left basin red and those that terminate in the right basin as blue. The two stable steady states of $F(x,v)$ show that our model predicts smaller cells to be bistable (Fig. 3), with streamlines similar to the bistability reported in [14] (reprinted in Fig. 3).
If we decrease the cell size significantly below $R_{0}=15{\mu\mathrm{m}}$, we find that the cells essentially migrate freely within the basins—very small cells rarely interact with the edges of the micropattern and almost never enter the bridge or successfully hop. Lastly, making cells larger than our default values does not affect their migration pattern: cells larger than the default $R_{0}=18{\mu\mathrm{m}}$ still limit-cycle with morphologies and streamlines that are highly similar to the default cell shown in Figures 3-3.
Why are smaller cells bistable in our model? When the cell is relatively small, its body does not overlap with the edges of the basin as much, so polarity is not as strongly suppressed near the micropattern edges. This leads to increased chances of developing a cell front anywhere along the cell periphery, not just in regions pointing at the bridge. This in turn increases the rate of repolarizations. Since smaller cells have lower chances of consistently polarizing into the bridge, they do not hop persistently, but rather stochastically, and are bistable. On the other hand, larger cells fill up the basin nearly entirely, leading to a consistent dampening of polarity along the cell periphery except around the regions pointing toward the bridge. This promotes persistent elongation into the bridge and thus limit cycles. The underlying mechanism is just that cells with larger spread area are forced to have more contact with the micropattern edges—so we expect that larger cells are more likely than small ones to limit-cycle in any model that drives cells to repolarize away from the edges of the micropattern. We expect this result would be fairly robust to changes in our detailed model assumptions.
Although the experiments of [14] do not monitor cell size quantitatively, their example movies show that MDA-MB-231 cells are typically spread over a larger area compared to MCF10A cells. There is also evidence that cancer cells have increased lamellipodial activity compared to their healthy counterparts [43]. Additionally, the absence of the tumor-suppressing micro RNA miR-200c has been shown to result in increased sizes of gastric cancer cells [44]. Collectively, these findings support the notion that cancer cells tend to have larger spread areas, rendering our model’s nomination of cell size as a predictor of motility behavior in two-state confinement plausible.
III.3 Cell tension could modulate bistability on two-state substrates
To hop between the basins, cells must elongate into the narrow bridge and tolerate a great deal of shape deformations (see the snapshots in Fig. 3). When $\gamma$, which sets the cell membrane tension, is low, i.e., the cell is soft, deformations are not very costly and can occur more easily. At high $\gamma$, the cell is stiff and shape deformations are costly. Therefore, we would expect soft cells to hop readily and stiff cells to resist hopping. To test this hypothesis, we first perturb our default cell by increasing its membrane tension to $\gamma=1.8\gamma_{0}$ and then simulate its migration on the two-state micropattern. Contrary to our intuition, we observe that the stiff cell still limit-cycles (Fig. 4), albeit with a small subset of the phase-space that originates and gets stuck oscillating inside the bridge (yellow region near $x=0,v=0$). When we look at the migration of stiff cells ($\gamma=1.8\gamma_{0}$) on two-state substrates, we see an unexpected behavior: stiff cells round up when passing through the bridge, which causes parts of the cell body to temporarily escape the micropattern (see Movie M4). Further increasing stiffness does not dampen limit-cycles—we simply observe a greater fraction of cells escaping the micropattern. Why does this happen? While higher tension naturally tends to make the cell round up, the active forces will still push the cell across the bridge. In the lower-$\gamma$ simulations (Fig. 3), the cell’s boundary contracts to within the micropattern because the polarity quickly becomes negative when the cell leaves the micropattern, leading to an inward-pointing force. However, once the cell boundary manages to leave the micropattern region, e.g. because the membrane tension resisting increasing the cell perimeter in the inward contraction is too large to overcome, no force is locally generated to bring it back (Eq. 5). As a result, the cell will round up and migrate across the bridge. We are not confident if this behavior is likely to be observed in experiment; it certainly would be strongly altered if we explicitly modeled an adhesive energy between the cell and the micropattern or altered our assumptions about where the cell could generate force.
Based on our analysis above, we think we can avoid the suspect behavior of the cell leaving the micropattern if we decrease the strength of active motility in some way. We can dampen the intensity of cell polarity by increasing the timescale $\tau_{f}$ over which new stochastic activity is added. We repeat the above numerical experiments with $\tau_{f}=3.6$s (compared to the default value $\tau_{f}=2.2s$). Under this alternate parameter regime, where cell polarity is less intense, we observe stiff cells ($\gamma=1.8\gamma_{0}$) to be bistable (Fig. 4) and softer cells ($\gamma=0.8\gamma_{0}$) to limit-cycle.
Figure 4: Motility patterns of soft ($\gamma=0.8\gamma_{0}$) and stiff ($\gamma=1.8\gamma_{0}$) cells on the two-state micropattern. 4 Under the default simulation parameters, stiff cells mostly exhibit limit-cycles, but 4 they can transition to bistable behavior when accompanied by a reduction in the intensity of cell polarity. Streamlines are colored by equilibrium point; $\times$ indicates equilibrium points where streamlines terminate. Solid black lines are separatrices—crossing this line changes the long-run behavior of the cell. See Section II.4 for details.
Our analysis shows that cell tension—or more generally, factors controlling cell deformability—can control whether the cell shows a bistable or limit cycle behavior. In [14], cancerous cells limit-cycled and healthy cells were bistable. Evidence suggests that cell stiffness is a potential biomarker for cell malignancy with cancer cells tending to be softer [45, 46, 47, 48]. Taken together, this lends confidence to our model’s nomination of cell stiffness as a potential predictor of bistability on the two-state micropattern.
III.4 Model recapitulates “negative friction” on two-state and normal friction on rectangular geometries
Following [14], we take a cut through the acceleration footprint near $x=0$, which denotes the symmetry axis of the micropattern, and ask how the center-of-mass acceleration at $x=0$ varies as a function of velocity, $F(x\to 0,v)$. For typical cell migration on a 2D substrate, which can often be captured by an Ornstein-Uhlenbeck-type model $\dot{\bm{v}}=-\xi\bm{v}+\eta(t)$ [49, 50, 13, 51], we would expect that the average acceleration $F(x\to 0,v)$ would be linear, i.e., $F(0,v)=-\xi v$, akin to a normal friction. Instead, our model predicts a nonlinear relationship for cells migrating on the two-state micropattern, with cells with small velocities tending to accelerate. This “negative friction” regime is consistent with the experimental measurements of [14] (Fig. 5). $F(x\rightarrow 0,v)$ is the average acceleration of cells whose center of mass is in the middle of the bridge. When the center of mass is in the bridge middle, the cell is in a dumbbell-like shape: its front is about to enter the new basin and the rear is about to retract (see Fig. 3). Both of these features mean that the cell will likely accelerate, as once the rear enters the bridge, it is not longer blocking forward motion, and once the front enters the new basin, it can generate protrusive force over a broader range. We thus think that the negative friction seen here requires the constriction of the cell as it enters the bridge.
This negative friction is entirely due to the narrow constriction in the two-state micropattern. For cells migrating on rectangular micropatterns or fully constriction-free substrates, we observe a nearly linear acceleration-velocity relationship, which is also consistent with the experiments of [14] (Fig. 5).
Figure 5: (Top) Simulation: Center-of-mass acceleration near the micropattern center exhibits a nonlinear (nearly linear) velocity dependence in two-state (rectangular) geometries. (Bottom) Experimental acceleration-velocity relationships for cell nucleus from [14], reprinted with permission.
III.5 Perturbing the polarization patch model
A key assumption of our model is characterizing the stochastic behavior of cell polarity by generating patches of activity every $\tau_{f}$ seconds, with a profile of $\beta(t)f\left(|\bm{r}-\bm{\theta}(t)|;\sigma\right)$, with amplitude $\beta$ and location $\bm{\theta}$ chosen randomly. In this section, we perturb this activity module away from its default state and gauge how sensitive our predictions are to its various components. Figure 6 outlines all the variations considered. As discussed below, our predictions are robust against certain aspects of the activity module, while other parameters are tightly constrained to recapitulate experimental observations.
Two-state Micropattern
Rectangular Micropattern
Figure 6: An overview of how various perturbations in the spatially-correlated activity patch we add to cell polarity affect the deterministic migration of the default cell on the two-state (left column) and rectangular (right column) micropatterns. In each column, the middle figure illustrates the behavior of the default cell, except in the last row, where the left figure represents this. The perturbations considered are in 6 the mean and 6 standard deviation of the magnitude of the noisy activity patch, 6 the spatial extent of the correlated activity patch, and 6 the frequency of adding noisy activity to cell polarity. Streamlines are colored by equilibrium point; $\times$ indicates equilibrium points where streamlines terminate. Solid black lines are separatrices—crossing this line changes the long-run behavior of the cell. See Section II.4 for details.
III.5.1 Perturbing $\beta$, the magnitude of noisy activity
$\beta(t)\sim\mathrm{Gaussian}(\mu_{\beta},\sigma_{\beta})$ is a random variable that sets the magnitude of the stochastic patches we add to cell polarity. We find that our model predicts limit-cycles on two-state and stationary behavior on rectangular micropatterns relatively robustly against variations in both $\mu_{\beta}$ (Fig. 6) and $\sigma_{\beta}$ (Fig. 6). However, there are a few interesting features that can appear within the larger limit-cycle dynamics. For the two cases where $\mu_{\beta}=1.4\mathrm{min}^{-1}$ and $\sigma_{\beta}=2.8\mathrm{min}^{-1}$, a new behavior emerges—in addition to the limit-cycles—that is confined to the interior of the bridge. This suggests that cells that happen to enter the bridge of the two-state micropattern with near-zero velocity may become trapped there, and either perform a small limit-cycle or reach a stationary state. These behaviors at the bridge are ones we are slightly skeptical about observing in experiments akin to [14], in part because our model does not have an explicit representation of the nucleus, while [14] use the cell’s nucleus to parameterize $x$. We would expect the nucleus to not be stably trapped exactly in the middle of the micropattern.
Lastly, we note that for $\sigma_{\beta}=6.6\mathrm{min}^{-1}$, the limit-cycles do not reach from the left basin to the right basin, but instead the streamlines loop around a narrower region. This corresponds to an oscillation of the cell center of mass around the bridge in our plot of $F(x,v)$. When we inspect our movies of cell trajectories, we see that in this case, cells quickly re-enter the bridge once they have entered a basin (see Movie M6).
We varied $\mu_{\beta}$ and $\sigma_{\beta}$ over a range set by monitoring the behavior of the cell. For values lower than $\mu_{\beta}=.63\mathrm{min}^{-1}$ and $\sigma_{\beta}=2.8\mathrm{min}^{-1}$, there isn’t a sufficient buildup of polarity, and we do not observe the cells to hop. Similarly, for higher values, we see many cell body ruptures as polarity becomes large enough to split one cell into many. The perturbations in this section show that our predictions are not overly sensitive to the magnitude of the activity patch so long as $\beta(t)$ is within a reasonable range.
III.5.2 Perturbing $\sigma$, the spatial extent of correlated activity
When we add a patch of activity to the cell, it is spread over a region $f\left(|\bm{r}-\bm{\theta}(t)|;\sigma\right)=e^{-|\bm{r}-\bm{\theta}(t)|^{2}/2\sigma^{2}}$—the parameter $\sigma$ sets the spatial extent over which one stochastic event affects the polarity.
Figure 6 summarizes the motility of cells with different values of $\sigma$. At the low value of $\sigma=3.4{\mu\mathrm{m}}$, we still observe the cells to limit-cycle on the two-state micropattern. However, a new behavior is observed on the rectangular geometry—cells are now bistable, settling deterministically into the left and right sides of the micropattern with two distinct terminal points. We do not have a strong intuition for why cells prefer the sides of the rectangular micropattern when their activity patches are tightly localized. At the large value of $\sigma=12{\mu\mathrm{m}}$, we recover the original stationary behavior on the rectangular geometry, but we observe cells to be mostly bistable on the two-state substrate (note the small fraction of lattice sites terminating near $(x=0,v=0)$ in Fig. 6). While we do not know with certainty why disperse patches lead to nearly bistable behavior in our model, we think it has to do with how spatially-extended patches activate a broader set of contour points. This, in turn, may increase the likelihood of activity occurring elsewhere on the cell boundary, thereby reducing the probability of persistent singly-polarized morphology. We conclude that $\sigma$ significantly impacts the migration of the cell in our model, leading to a tight range of values within which we can recapitulate experimental observations.
III.5.3 Perturbing $\tau_{f}$, the frequency of noisy activity
The last parameter in our stochastic patch model is $\tau_{f}$, the frequency with which we add noisy patches to cell polarity. We find that our model robustly predicts limit-cycles on two-state micropatterns and stationary behavior on rectangular micropatterns when changes in this timescale are small (Fig. 6). When we nearly double $\tau_{f}$ from its default value, we observe limit-cycles not only on the two-state substrate but also on the rectangular micropattern. This may occur because as $\tau_{f}$ increases, the time between stochastic protrusion events becomes large enough for the cell to develop a robust singly-polarized state, which then leads to a cell migrating persistently along the rectangular micropattern. This result shows that even the seemingly simple and intuitive result that cells on the rectangular pattern have near-stationary dynamics can be constraining to a detailed model.
III.6 Can we simplify our model assumptions?
Having confirmed our model’s ability to recapitulate the findings of [14], we now investigate whether a simplified version of it can achieve the same success. We do this by knocking out various aspects of the model and observing how each affects the migration dynamics of our default cell on the two-state and rectangular micropatterns.
We begin by asking if the two mechanisms by which the cell integrates information about substrate geometry, i.e., the filopodial probability in Eq. 7, and the coupling to $\chi$ in Eq. 6, are both needed. We knock out the filopodial sensing mechanism by simplifying the probability of Eq. (7) to $p(\bm{\theta})\sim\mathbb{P}(\bm{\theta})$. Upon this knockout, we observe no significant change in the motility pattern or the velocity dynamics of the default cell on either the two-state or rectangular micropatterns (Fig. S2). However, we observe a notable change in the distribution of “hopping times”, defined as the duration of time spent in a basin of the two-state geometry before hopping (Fig. 7). While our default model predicts a tightly-peaked distribution, consistent with the findings of [14], this perturbation leads to a broader density. The broadening is expected; without the filopodial probability, cells have reduced ability to aim their protrusive activity toward the bridge, which naturally delays the hopping process and extends the hopping time. Considering the hopping time distribution, we opt to retain the filopodial probability to ensure a more faithful reproduction of the observations of [14].
Figure 7: Distribution of “hopping times” predicted by the default model vs. one with the filopodial mechanism absent.
Next we knock out the coupling of cell polarity to the substrate by taking $\tau_{\chi}$ to infinity (Fig. 8). In this case, the cell is still largely confined within the substrate due to the filopodial sensing mechanism disfavoring any polarization that would cause the cell to escape. While motility in the rectangular substrate remains unaffected (Fig. S3), migration in the two-state micropattern develops multiple equilibrium points: while the two points at either basin are robust to sampling uncertainties, the putative equilibrium points near the origin are not (see Fig. S12). We thus interpret the streamlines of Fig. 8 as effectively suggesting a bistable behavior. Additionally, the acceleration at $x\to 0$ no longer displays the pronounced negative friction ($F(x\to 0,v)\sim+v$) that is characteristic of migration in the two-state micropattern (compare Fig. 8 to the blue and green curves in Fig. 5). Both the filopodial probability and the cell-micropattern coupling sense spatial cues, altering polarity in response to the presence of the micropattern, yet only the latter disrupts the motility dynamics of the cell. Why is that? The key difference between each sensory input is likely how they affects cell polarity. The filopodial sensing mechanism changes the distribution of polarity activation patches. Because the average effect of these activation patches is positive, redistributing these activity patches will not generally create a negative polarity at any point in the cell. By contrast, the micropattern coupling term $-\tau_{\chi}^{-1}\chi\phi$ can make the polarity negative and lead to the membrane retracting away from the micropattern edge. When the cell is about to enter the bridge and the cell-micropattern coupling is present, the constant decrease in polarity pinches the membrane near the corners of the constriction and facilitates the formation of tubular protrusive fronts while holding back the cell body (Fig. 3). Without this coupling, the cell does not contract its membrane at the pinch points of the constriction; it fails to grow tubular fronts and adopts a morphology that is unfazed by the micropattern constriction (see Fig. 8 and Movie M7). Considering this morphological change, the absence of a clear negative friction dynamic with this knockout is not surprising given we think that squeezing through the bridge is necessary for exhibiting nonlinear dynamics (see our argument in Section III.4).
Figure 8: 8 Sample simulation snapshots of our default cell migrating on the two-state micropattern, along with 8 a summary of its motility dynamics when cell-confinement coupling is turned off, i.e., $\tau_{\chi}\rightarrow\infty$.
Lastly, we knock out the perimeter-induced degradation of cell polarity by taking $s_{0}\rightarrow\infty$ and observe no significant change in the motility pattern or the velocity dynamics of the default cell on either micropatterned geometry (Fig. S4). Although we could safely eliminate this assumption and still recapitulate the experimental observations, we choose to keep it because it leads to a cell morphology that is closer to that of MDA-MB-231 cells in [14]. Without it, there is a higher chance for cells to become bipolar midway through the bridge or for one bipolar cell to break apart into two (see Movie M8). Given the clear experimental rationale for this sort of term [37], we keep the perimeter-dependent global inhibition in our default model.
IV Discussion
Brückner et. al. revealed the distinct migratory dynamics of cancerous MDA-MB-231 and healthy MCF10A cells crawling on two-state and rectangular micropatterned substrates by inferring the cell’s equation of motion from observations. Here, we have proposed a single model of the cell and showcased its successful recapitulation of observations in [14]. Our model correctly predicts limit cycle, stationary, and bistable behavior depending on cell properties and the geometry of the micropattern. It also successfully predicts the nonlinear (nearly linear) dynamics of cells on the two-state (rectangular) substrate, consistent with [14] (Fig. 5).
The behavior of MDA-MB-231 cells on rectangular and two-state micropatterns is radically different, with cells switching from a weakly-polarized fluctuating state to a highly persistent oscillation (Figs. 3 and 3). We have been able to recapitulate this shape sensing with a very minimal assumption—that the cell’s polarity is locally inhibited by absence of cell-fibronectin contact. This assumption seems well-supported by existing literature showing strong effects of fibronectin on Rho GTPase polarity [36]. Our simple assumption may provide a mechanistic view of what could drive the geometrically-adaptable polarity hypothesized by [52].
Within our model, we can switch cells between showing limit-cycle dynamics and bistability either by changing physical parameters like cell size or tension, or by changing polarity dynamics like the size of protrusion fluctuations. Our results indicate that these distinct motility patterns can reflect quantitative changes in the cell, rather than requiring fundamentally different assumptions in how we model MCF10A and MDA-MB-231 cells. Consistent with this idea, it was later found that subpopulations of cancerous cells, derived from their inherent cell-to-cell variability, can also exhibit distinct motility behaviors [53]. Additionally, the expression of the tumor-suppressing micro RNA miR-200c in MDA-MB-231 cells has recently been shown to induce a switch in their motility behavior on the two-state micropattern, with cells expressing miR-200c becoming bistable [54]. While miR-200c has many targets and its impact on cell migration is complex, there is evidence that its expression can lead to increased cell stiffness [55, 56] and decreased cell size [44]. Taken together, these findings complement our predictions that dramatic, qualitative changes in cell behavior can possibly be explained by simple, quantitative changes in the cell.
If the key difference between MCF10A and MDA-MB-231 cells were the cell spread area, which we have shown is sufficient to explain a transition between limit cycle and bistable behaviors, this could be easily tested through both quantification of cell spread areas and variation of the micropattern size. Within our model, simulated cells with similar $A_{\textrm{cell}}/A_{\textrm{basin}}$ have similar dynamics, assuming all other parameters are held constant. To establish further confidence in our findings and probe the limits of our model, we would explore motility in other geometries. For example, there are experimental observations for MDA-MB-231 cells on spiked ratchets [3], as well as various two-state geometries with circular-square and circular-triangular basins [57], all of which provide rich context for further comparisons. Further experimental studies will also be necessary to validate our suggestion of cell properties as predictors of cell motility.
In our application of the data-driven approach of [14], we discovered that uncertainties in the inferred nonlinear dynamics—driven by finite sampling—can be significant. This is despite our large dataset of 4800 simulations spanning for 48 hours, providing roughly 24x as much trajectory time as observed experimentally. This means a single streamline plot should be taken with some skepticism. For instance, the bistable dynamics seen in Fig. 6 with $\sigma=12{\mu\mathrm{m}}$ generates a streamline plot that is slightly asymmetric at the center, and has a small domain of attraction at the center. These features, though, are not robust to sampling (see Appendix E). We suggest that the bootstrap approach we take in Appendix E should be used as an essential robustness check.
Acknowledgements.
PZ and BAC are supported by NIH R35 GM142847. This work was carried out at the Advanced Research Computing at Hopkins (ARCH) core facility (rockfish.jhu.edu), which is supported by the National Science Foundation (NSF) grant number OAC 1920103. We thank Kurmanbek Kaiyrbekov and Yongtian Luo for a close reading of a draft of this paper.
Appendix A Numerically evolving the phase field equations of motion
The phase field equations of motion are
[…]
where the total free energy of the cell is $\mathcal{F}=\mathcal{F}_{\mathrm{CH}}+\mathcal{F}_{\mathrm{area}}$. Given the definitions in Eqs. (2) and (3), the functional derivatives are
$\frac{\delta\mathcal{F}_{\rm CH}}{\delta\phi}=36\frac{\gamma}{\lambda}\phi(2\phi-1)(\phi-1)-2\lambda\gamma\nabla^{2}\phi,$ (10)
and
[…]
With $\frac{\delta\mathcal{F}}{\delta\phi}$ worked out, we numerically evolve Eq. (9) on a 200$\times$200 grid with a spatial resolution of $\Delta x=1.5{\mu\mathrm{m}}$ and simulation timestep of $\Delta t=0.36$s [26]. We also leverage scikit-image’s measure function [58] to find contour points at the 1/2 level and matplotlib’s streamplot [59] to trace out streamlines.
Appendix B Constructing acceleration footprints from simulated trajectory data
We record the $x$ position of cell center-of-mass every $\Delta t=3$ min in our simulations, and at the end of each run, we calculate the velocity and acceleration from this time series:
$v(t)$ $=\frac{x(t+\Delta t)-x(t)}{\Delta t},$ (12)
$a(t)$ $=\frac{x(t+\Delta t)-2x(t)+x(t-\Delta t)}{\Delta t^{2}}.$ (13)
Note that $\Delta t$ here is not the time step of the previous section, but the time resolution with which we measure the cell’s center of mass. With the above equations, we compute the tuples $(x,v,a)$ for all center-of-mass trajectories observed across all 4800 simulations, which amounts to roughly 2.3M records. We construct the $xv$ phase-space by discretizing the domain of each coordinate with $n=32$ bins, where each bin has dimensions $dx=(x_{\rm max}-x_{\rm min})/n$ and $dv=(v_{\rm max}-v_{\rm min})/n$. Considering the spatial and velocity domains typical to our simulations, our bins are approximately $3{\mu\mathrm{m}}\times 11{\mu\mathrm{m}}$/hr. To populate this phase-space, we group together the set of all tuples $(x,v,a)$ that fall into the $(x,v)$ bin and assign that bin the average acceleration of the set. Repeating this procedure for all the $n\times n$ bins present yields the acceleration footprint $F(x,v)=\langle a(x,v)|x,v\rangle$.
Appendix C A dynamically-evolving cell boundary is required for bipolar morphology
In Section III.1, we outlined our model’s ability to recapitulate limit-cycles on two-state and stationary behavior on rectangular substrates. The emergence of persistently hopping cells is plausible considering our assumption of positive feedback for cell polarity (see Eq. 6)—this mechanism amplifies small differences in polarity into one hot spot of activity, thus promoting singly-polarized cells that can migrate persistently. The same principle underlies the migration of the cell on the rectangular micropattern, yet the cell adopts a bipolar morphology on average, leading to near zero speeds and the observed stationary states (compare Movies M1 and M2). To investigate how our assumptions for cell polarity can lead to bipolar cells on the rectangular micropattern, we decided to study the evolution of cell polarity in isolation. This simplified version of the model represented the cell with a static, circular phase field initialized in the center of the rectangular geometry. Here, we did not update the cell interface itself, i.e., $\frac{\partial}{\partial t}{\phi}_{\rm static}=0$, but did evolve the polarity according to Eqs. (6) and (7) on the static cell boundary. Surprisingly, we observed the stochastic evolution of cell polarity to always localize into a single hot spot of activity, consistently yielding a singly-polarized cell (Fig. S1). Since the only change we made was making the phase field static, we conclude that the emergence of bipolar cell morphologies in our model requires the cell interface to be dynamic.
Figure S1: A sample of the steady-state distribution of cell polarity as evolved according to Eqs. (6) and (7) on a static phase field initialized on the rectangular micropattern. Although each sample starts with a random initial distribution and evolves stochastically, all of them equilibrate to a singly-polarized configuration.
Appendix D Predicted motility dynamics in response to simplifying model knockouts
Figure S2: Predicted motility dynamics of our default cell on S2 the two-state and S2 the rectangular micropatterns with the filopodial mechanism absent, i.e., $p(\bm{\theta})=\mathbb{P}(\bm{\theta})$.
Figure S3: Predicted motility dynamics of our default cell on the rectangular micropattern with the degradation of polarity outside the micropattern absent, i.e., $\tau_{X}\to\infty$.
Figure S4: Predicted motility dynamics of our default cell on S4 the two-state and S4 the rectangular micropatterns with the perimeter-induced degradation of polarity absent, i.e., $s_{0}\to\infty$.
Appendix E Showcasing the uncertainty of phase-space separatrices
Our deterministic acceleration footprints $F(x,v)$ have associated uncertainties since we infer them from a finite number of simulations. Here, we focus on footprints that exhibit multiple phases and seek to demonstrate the typical variability we can expect when interpreting the motility patterns encoded by them. Note that we specifically highlight the uncertainty of these acceleration footprints because we characterize them by their separatrices and want to better understand the stability of these boundaries. To assess variability, we employ bootstrapping on our original set of simulations. Each bootstrap involves sampling 4800 times from the set of independent phase field simulations with replacement, allowing us to infer $F_{k}(x,v)$ for each bootstrapped ensemble $k$. We generate a total of 48 such bootstrapped ensembles, i.e., $k=[1,2,...,48]$. The below figures showcase the variability of their respective acceleration footprint, each consisting of three panels:
  • (a)
    This is the original $F(x,v)$ inferred from our phase field simulations. A copy of this is in the main text.
  • (b)
    This is a heatmap showing the number of distinct categories each phase-space lattice site has ever been a member of across the entire set of bootstrapped ensembles. There is a maximum of 3 phases a lattice site can be classified as based on whether it terminates at the left, center, or right of the micropattern. In this view, lattice sites that fluctuate are colored darker and serve as a visual representation of the variability we can expect when delineating the boundary of distinct phases.
  • (c)
    This is an exemplary view of 9 out of the 48 bootstrapped acceleration maps. It is a useful way to explicitly demonstrate the variability we can expect in separatrix boundaries.
Refer to caption (a) Refer to caption (b)
Refer to caption (c)
Figure S5: Overview of variability in motility for the cell-micropattern pair presented in Fig. 3.
Refer to caption (a) Refer to caption (b)
Refer to caption (c)
Figure S6: Overview of variability in motility for the cell-micropattern pair presented in Fig. 4.
Refer to caption (a) Refer to caption (b)
Refer to caption (c)
Figure S7: Overview of variability in motility for the cell-micropattern pair presented in Fig. 4.
Refer to caption (a) Refer to caption (b)
Refer to caption (c)
Figure S8: Overview of variability in motility for $\mu_{\beta}=1.4\mathrm{min}^{-1}$ perturbation on the two-state geometry presented in Fig. 6.
Refer to caption (a) Refer to caption (b)
Refer to caption (c)
Figure S9: Overview of variability in motility for $\sigma_{\beta}=2.8\mathrm{min}^{-1}$ perturbation on the two-state geometry presented in Fig. 6.
Refer to caption (a) Refer to caption (b)
Refer to caption (c)
Figure S10: Overview of variability in motility for $\sigma=12{\mu\mathrm{m}}$ perturbation on the two-state geometry presented in Fig. 6.
Refer to caption (a) Refer to caption (b)
Refer to caption (c)
Figure S11: Overview of variability in motility for $\sigma=3.4{\mu\mathrm{m}}$ perturbation on the rectangular geometry presented in Fig. 6.
Refer to caption (a) Refer to caption (b)
Refer to caption (c)
Figure S12: Overview of variability in motility for the cell-micropattern pair presented in Fig. 8.
Appendix F Legend for the attached simulation movies
All simulations are run for 24 hours.
  1. 1.
    Movie 1: Default cell limit cycling on the two-state micropattern
  2. 2.
    Movie 2: Default cell migrating on the rectangular micropattern
  3. 3.
    Movie 3: Default, small cell with $R_{0}=15{\mu\mathrm{m}}$ migrating on the two-state micropattern
  4. 4.
    Movie 4: Default, stiff cell with $\gamma=1.8\gamma_{0}$ migrating on the two-state micropattern
  5. 5.
    Movie 5: Default cell freely migrating on an unconstrained micropattern (axes are in ${\mu\mathrm{m}}$)
  6. 6.
    Movie 6: Default, noisy cell with $\sigma_{\beta}=6.6\mathrm{min}^{-1}$ migrating on the two-state micropattern
  7. 7.
    Movie 7: Default cell with $\tau_{\chi}\to\infty$ migrating on the two-state micropattern
  8. 8.
    Movie 8: Default cell with $s_{0}\to\infty$ migrating on the two-state micropattern